# MoPE: Parameter-Efficient and Scalable Multimodal Fusion via Mixture of Prompt Experts

### v2

@@ Abstract
Despite the demonstrated parameter efficiency of prompt-based multimodal fusion methods, their limited adaptivity and expressiveness often result in suboptimal performance compared to other tuning approaches. In this paper, we address these limitations by decomposing the vanilla prompts to adaptively capture instance-level features. Building upon this decomposition, we introduce the mixture of prompt experts (MoPE) technique to enhance the expressiveness of prompt tuning. MoPE leverages multimodal pairing priors to route the most effective prompt on a per-instance basis. Compared to vanilla prompting, our MoPE-based fusion method exhibits greater expressiveness, scaling more effectively with the training data and the overall number of trainable parameters. We also investigate regularization terms for expert routing, which lead to emergent expert specialization during training, paving the way for interpretable soft prompting. Extensive experiments across six multimodal datasets spanning four modalities demonstrate that our method achieves state-of-the-art results for prompt fusion, matching or even surpassing the performance of fine-tuning while requiring only 0.8% of the trainable parameters.
Introduction
@@
Unimodal pre-trained models like BERT (Radford et al. 2019) and Swin Transformer (Liu et al. 2021b) have excelled in transferring capabilities to a wide range of tasks. By comparison, extending the pretraining-finetuning paradigm to multimodal applications involves additional complexities, particularly regarding flexibility and cost. First, the modality-paired pretraining can be inflexible due to its dependence on specific configurations. This can limit flexibility when there is a scarcity of certain modality pairs, or when new architectures emerge, necessitating resource-intensive retraining. In contrast, unimodal pretrained models are more readily available and offer greater flexibility, as they can be independently updated and combined for multimodal tasks. Second, as these foundation models scale in size, fine-tuning them becomes increasingly expensive, which further restricts their application to downstream tasks. To democratize foundation models, a compelling question arises: How can we efficiently combine separately pretrained unimodal models for multimodal tasks?
@@
Figure 1: High-level motivation of MoPE-based multimodal fusion. (a) Vanilla prompt tuning learns a globally shared long prompt for all instances, which may not be optimal for each instance. (b) MoPE decomposes the global prompt into multiple short and specialized prompt experts to improve its adaptivity and expressiveness.
@@
While originally proposed for transfer learning (Li and Liang 2021; Jia et al. 2022; Zhou et al. 2022b), recent research has revealed that prompt tuning could also be a parameter-efficient paradigm for multimodal fusion (i.e., prompt fusion). Typically, this is achieved by representing one modality with several prompts, which are fed into the frozen Transformer of another modality to achieve fusion (Tsimpoukelli et al. 2021; Liang, Zhao, and Schütze 2022; Li et al. 2023; Yang et al. 2022b). However, directly using prompts for fusion could yield suboptimal results, particularly due to their limited adaptivity and scalability. Notably, while prompt tuning generally performs well in low-data regimes, it can be less effective when applied to full-shot training on the entire dataset with a challenging objective (Gao et al. 2022; Liang, Zhao, and Schütze 2022; Tsimpoukelli et al. 2021; Yang et al. 2022a; Lester, Al-Rfou, and Constant 2021). This reduced efficacy could be attributed to at least two factors. First, the vanilla prompt was originally developed for unimodal learning, which optimizes globally-shared prompts for all instances (Jia et al. 2022; Li and Liang 2021; Tsimpoukelli et al. 2021). When directly applied for prompt fusion (Tsimpoukelli et al. 2021; Li et al. 2023; Liang, Zhao, and Schütze 2022), it fails to adaptively exploit the multimodal interplay and can lead to sub-optimal representations for each input instance. Second, the relatively limited representation power (compared to fine-tuning) can lead to underfitting in multimodal datasets with a long-tail distribution and complex cross-modal mapping (Liu et al. 2021a; Petrov, Torr, and Bibi 2023; Wang et al. 2023).
@@
To address these challenges, increasing the number of learnable prompts, known as “length-scaling”, appears to be a straightforward solution. However, numerous studies have shown that the performance gains from length-scaling quickly reach saturation in both transfer-learning (Gao et al. 2022; Lester, Al-Rfou, and Constant 2021) and fusion settings (Liang, Zhao, and Schütze 2022; Tsimpoukelli et al. 2021; Yang et al. 2022a). Furthermore, over-length prompts may even lead to worse results (Jia et al. 2022; Yang et al. 2022a; Khattak et al. 2023; Li and Liang 2021; Kim et al. 2023). Recent theoretical analyses have substantiated these empirical observations, highlighting the limited expressiveness of prompt tuning (Wang et al. 2023; Petrov, Torr, and Bibi 2023). Overall, evidence from both empirical and theoretical analyses reveals that simply increasing prompt length is insufficient to address the limitations of prompt fusion.
To improve the adaptivity and expressiveness of prompt fusion, we propose decomposing the unified, long prompt into multiple short prompts with specialized responsibilities, as depicted in Fig. 1. We introduce an instance-wise adaptive prompt decomposition that separates the vanilla prompt into static, dynamic, and mapped prompts to adaptively capture global and local features. The foundation of our dynamic prompt generation lies in the proposed MoPE (Mixture of Prompt Experts) design, where we maintain a fixed prompt length in the self-attention and enhance expressiveness by learning multiple prompt experts and a multimodal router to select the most suitable prompt for each instance.
@@
We systematically compare our MoPE-based fusion with previous prompt fusion methods on a total of six datasets, and demonstrate consistently superior performance with higher parameter-efficiency. Furthermore, we showcase the superior scalability of our method both in terms of parameter and data. Our analysis indicates that increasing the number of experts (i.e., “expert-scaling”) is more scalable than length-scaling, with monotonic performance gains and the avoidance of performance deterioration with an overly-long prompt. Our method also scales better with more training data. By introducing regularization terms for expert routing, we observe the emergence of specialized prompt experts after end-to-end training, resulting in high interpretability. Our key contributions are summarized as follows:
  • We propose an instance-wise adaptive prompt decomposition, which decomposes the vanilla prompt into static, dynamic and mapped prompt.
@@
  • We introduce the MoPE technique for instance-wise dynamic prompt generation, which scales up the expressiveness of prompt tuning for multimodal fusion.
@@
  • Extensive experiments on six datasets spanning four modalities demonstrate state-of-the-art performance and parameter efficiency for prompt-based multimodal fusion.
Related Works
Prompt Tuning for Transfer Learning. Prompt tuning (Li and Liang 2021; Lester, Al-Rfou, and Constant 2021) learns continuous token embeddings as additional input to a frozen pretrained model for transfer learning of Transformer-based models. It is widely used in various modalities (Jia et al. 2022; Bahng et al. 2022; Liu et al. 2022; Lin et al. 2023; Jiang, Liu, and Chen 2023; Khattak et al. 2023; Zhou et al. 2022a). A common observation is good transfer learning performance in low-shot scenarios, yet its performance is less comparable to fine-tuning when abundant training instances are available (Li and Liang 2021; Gao et al. 2022). Moreover, increasing prompt length quickly reaches performance saturation, and over-length prompts might lead to worse results (Yang et al. 2022a; Jia et al. 2022; Kim et al. 2023). Recent theoretical analyses (Petrov, Torr, and Bibi 2023; Wang et al. 2023) reveal that the expressiveness of prompt tuning is lower than that of fine-tuning. In this work, we tackle this challenge by scaling up the expressiveness of prompts with a Mixture of Experts (MoE)-like design.
Multimodal Fusion via Prompt. Apart from model adaptation, prompts can also be used to fuse separately pretrained unimodal models for multimodal tasks. Frozen (Tsimpoukelli et al. 2021) first introduced a method where the visual representation is mapped as a few input tokens to query frozen language models (LMs). PromptFuse and BlindPrompt (Liang, Zhao, and Schütze 2022) improved upon this by introducing tunable prompts to the LM for cross-modal alignment. PICa (Yang et al. 2022b) translates images into discrete text captions to prompt frozen LMs. These methods treat tokens from different modalities independently and lack cross-modal interaction. Recently, PMF (Li et al. 2023) introduced an interactive prompt fusion method for the vision and text modalities, building on the strong assumption that both encoders are white-box and have the same Transformer architecture. In this paper, we investigate how to adaptively interact with two or more modalities with minimal assumptions about the model architecture.
MoE in Transformers. MoE is a computationally efficient technique to scale up models, including Transformers (Lepikhin et al. 2020; Riquelme et al. 2021; Mustafa et al. 2022; Fedus, Zoph, and Shazeer 2022). The fundamental approach involves inserting MoE layers, usually composed of multiple feed-forward networks (FFNs) acting as experts, into the standard Transformer architecture. A router is learned to route each token embedding to the most suitable expert(s) for reducing computational overhead (Qin and Eisner 2021; Riquelme et al. 2021; Shazeer et al. 2017). In this paper, we do not focus on reducing computational cost; instead, we introduce the MoE design into prompt fusion to scale up its adaptiveness and expressiveness.
Method
@@
Preliminary: Vanilla Prompt Tuning
Consider a Transformer (Vaswani et al. 2017) or its variants used to extract features from an embedded input sequence $\mathbf{x}^{0}\in\mathbb{R}_{s\times d_{x}}$, where $s$ is the sequence length and $d_{x}$ is the embedding dimension. Prompt-tuning freezes all pre-trained parameters and optimize a small amount of continuous embeddings (i.e., prompts) $\mathbf{P}\in\mathbb{R}_{l\times d}$ concatenated to the input of each layer, where $l$ is length of the prompt. The input of the $i$-th layer layer $L^{i}$ could be denoted as:
@@
Succinctly, prompt tuning works by biasing the pretrained attention pattern in each Transformer layers (Petrov, Torr, and Bibi 2023). As opposed to fine-tuning, this biasing has strictly limited expressiveness in theory (Petrov, Torr, and Bibi 2023; Wang et al. 2023), meaning that there are tasks that are un-learnable even with $l\to\infty$. The empirical expressiveness of vanilla prompting is even lower than its theoretical upper bound, due to competing optimization when $l>1$ (Petrov, Torr, and Bibi 2023). In other words, although longer prompt gives more trainable parameters, finding such a long prompt becomes challenging, and a suboptimal long prompt can lead to performance degradation instead. In this paper, we do not aim to increase the theoretical upper bound of prompt-tuning; instead, we approach the problem by narrowing the gap between its empirical and theoretical expressiveness.
Instance-wise Adaptive Prompt Decomposition
Previous prompt fusion methods (Liang, Zhao, and Schütze 2022; Yang et al. 2022b) optimize a global prompt that is shared across all instances, neglecting the interplay between different modalities and instance-specific features. With paired multimodal input as a prior, we aim to instance-wisely condition the prompting of one modality on the other(s) for better adaptivity. To achieve this goal, we first adopt a sequential fusion pipeline. Let $x\in\mathbb{X},y\in\mathbb{Y}$ be a pair of multimodal inputs, and $\mathcal{E}_{\mathbb{X}}$, $\mathcal{E}_{\mathbb{Y}}$ be the encoder of each modality. Depending on the task intrinsic, we assign a fusion direction where the encoding of main modality $\mathbb{X}$ is conditioned on representation of complementary modality(ies) $\mathbb{Y}$.
@@
Mixture of Prompt Experts
@@
where $\tau=0.1$ is the temperature hyper-parameter, and $\epsilon$ is sampled noise (Riquelme et al. 2021). When there are more than one complementary modality, the additional embedding could be easily extended by learning additional projections. The dynamic prompt is obtained by a convex combination of all experts at this layer according to the routing score:
@@
Regularizing Expert Routing
We empirically find that a few experts tend to dominate across all instances during training, a phenomenon that has also been observed in previous MoE methods (Eigen, Ranzato, and Sutskever 2013; Shazeer et al. 2017). In this section, we introduce methods to circumvent degenerated routing and aid expert specialization.
Orthogonal Routing Embedding. To avoid the expert being selected in initialization being exposed to a larger gradient signal, we freeze the routing embeddings $\{\mathbf{k}\}_{j=1}^{k}$ to suppress self-reinforcing expert domination. Our finding indicate that an orthogonal initialized (Saxe, McClelland, and Ganguli 2013) and non-learnable routing embedding outperform learned one, while requiring fewer trainable parameters.
Importance Loss. To further aid expert specialization, we add an additional importance loss (Shazeer et al. 2017; Riquelme et al. 2021) to penalize dominant experts. For a batch of input $\mathbf{B}=\{(x_{0},y_{0}),(x_{1},y_{1}),...,(x_{b},y_{b})\}$, the importance of expert-$j$ is defined as the summed routing score in this batch, $\operatorname{Imp}(\mathbf{E}_{j})=\sum_{(x,y)\in\mathbf{B}}\mathbf{r}_{j}$. The importance loss is calculated as the mean coefficient of variation of all experts’ importance averaged across all layers:
@@
where $\operatorname{\sigma}(\cdot)$ is the stop-gradient operator to prevent error propagation of this loss term when the coefficient of variation is less than a pre-defined threshold $\gamma=0.1$. While this loss was initially proposed for balancing computational budgets (Shazeer et al. 2017; Riquelme et al. 2021), we adapt its use for promoting expert specialization. We add an additional threshold constraint due to our instance-wise routing, which is more likely to have a larger coefficient of variation than per-token routing.
@@
Acc ($\uparrow$) Acc ($\uparrow$) F1-macro/F1-micro ($\uparrow$) Pre ($\uparrow$) Pre ($\uparrow$)
@@
Experiments
Datasets. Following previous prompt fusion methods (Liang, Zhao, and Schütze 2022; Li et al. 2023), we conduct experiments on MM-IMDB (Arevalo et al. 2017) and UPMC Food-101 (Wang et al. 2015) for classification, SNLI-VE (Xie et al. 2019) for visual entailment (with image premise only), MUsTARD (Castro et al. 2019) for sarcasm detection in speaker-dependent setting. We additionally test our method on the RefCOCO and RefCOCO+ (Kazemzadeh et al. 2014) dataset for referring expression segmentation.
Architecture Details. In main experiment, we use Swin-B (Liu et al. 2021b) as the image encoder, Bert-base (Devlin et al. 2018) as the text encoder, Wav2Vec2 (Baevski et al. 2020) as the audio encoder. For the video modality, we use the same frame averaging technique by adapting a ViT (Dosovitskiy et al. 2020) encoder, which follows PromptFuse (Liang, Zhao, and Schütze 2022). Image is used as main modality unless otherwise specified. Following the experiment setup in (Jia et al. 2022; Li et al. 2023; Liang, Zhao, and Schütze 2022), we finetune dataset-specific heads. We implement the mapper as a two-layer MLP with GeLU nonlinearily. Regarding prompt, we set $l=6$ and use $k=16$ experts by default, and the prompts are applied to all layers of the main modality encoder. Vanilla prompt tuning (Li and Liang 2021) with $l^{\prime}=4$ is used to tune the $\mathcal{E}_{\mathbb{Y}}$.
Training Details. All models are trained for 20 epoches, using the AdamW (Loshchilov and Hutter 2017) optimizer with a learning rate of $4\times 10^{-4}$ for main modality and $5\times 10^{-4}$ for complementary modality. All models are trained with a RTX-4090 GPU.
Further experiment details could be found in the Supplementary Material.

### v3

# MoPE: Mixture of Prompt Experts for Parameter-Efficient and Scalable Multimodal Fusion

@@ Abstract
Despite the demonstrated parameter efficiency of prompt-based multimodal fusion methods, their limited adaptivity and expressiveness often result in suboptimal performance compared to other tuning approaches. In this paper, we introduce the Mixture of Prompt Experts (MoPE), the first technique designed to overcome these limitations by decomposing standard prompts to capture instance-level features adaptively. Building on this decomposition, MoPE enhances prompt fusion’s expressiveness by leveraging multimodal pairing priors to route the most effective prompt for each instance dynamically. Compared to vanilla prompting, our MoPE-based fusion method exhibits greater expressiveness, scaling more effectively with the training data and the overall number of trainable parameters. We also investigate regularization terms for expert routing, which lead to emergent expert specialization with enhanced adaptiveness and interpretablity. Extensive experiments across six multimodal datasets spanning four modalities demonstrate state-of-the-art performance for prompt fusion, matching or even surpassing the performance of fine-tuning while requiring only 0.8% of the trainable parameters. Project homepage: https://github.com/songrise/MoPE.
1 Introduction
@@
Unimodal pre-trained models like Bert [30] and Swin Transformer [25] have excelled in transferring capabilities to a wide range of tasks. By comparison, extending the pretraining-finetuning paradigm to multimodal applications involves additional complexities, particularly regarding flexibility and cost. First, the modality-paired pretraining can be inflexible due to its dependence on specific configurations. When there is a scarcity of certain modality pairs, or when new architectures emerge, it necessitates resource-intensive retraining. In contrast, unimodal pretrained models are more readily available and offer greater flexibility, as they can be independently updated and combined for multimodal tasks. Second, as these foundation models scale in size, fine-tuning them becomes increasingly expensive, which further restricts their application to downstream tasks. To democratize foundation models, a compelling question arises:
How can we efficiently combine separately pretrained unimodal models for multimodal tasks
@@
Figure 1: High-level motivation of MoPE-based multimodal fusion. (a) Vanilla prompt tuning learns a globally shared long prompt for all instances, which may not be optimal for each instance. (b) MoPE decomposes the unified long prompt into multiple specialized and short prompt experts to improve its adaptivity and expressiveness.
@@
While originally proposed for transfer learning [19, 11, 43], recent research has revealed that prompt tuning could adapted for parameter-efficient multimodal fusion (i.e., prompt fusion). Typically, this is achieved by representing one modality with several prompts and feed into the frozen Transformer of another modality [34, 21, 20, 41]. However, directly using prompts for fusion could yield suboptimal results. It has been observed in various papers that while prompt tuning generally performs well in low-data regimes, it can be less effective when applied to full-shot training on larger datasets with a challenging objective [9, 21, 34, 40, 18].
The reduced efficacy of prompt fusion could be attributed to its limited adaptivity and expressiveness. Specifically, these methods typically employ globally shared prompts [34, 20, 21] adopted from vanilla prompt tuning for all instances, which fails to capture the nuanced, instance-specific multimodal interactions. Additionally, the constrained expressiveness of prompt (compared to fine-tuning) can lead to underfitting in multimodal datasets with a long-tail distribution and complex cross-modal mapping [24, 28, 37].
@@
To address these challenges, increasing the number of learnable prompts, known as “length-scaling”, appears to be a straightforward solution. Nevertheless, the performance gains from length-scaling quickly reach saturation in both transfer-learning [9, 18] and fusion settings [21, 34, 40]. Furthermore, over-length prompts may even lead to worse results [11, 40, 14, 19, 16]. Recent theoretical analyses have substantiated these empirical observations [37, 28], particularly regarding the difficulty in optimizing a unified long prompt vector.
In this work, we present the first approach that specifically targets the enhancement of both adaptivity and expressiveness in prompt fusion. Our high-level idea is to divide-and-conquer the problem space utilizing multimodal pairing information, as depicted in Fig. 1. Specifically, we propose decomposing the unified, long prompt into three types of specialized, short prompts that are instance-wisely adaptive. To synthesize the most effective prompt per-instance, the MoPE module (Mixture of Prompt Experts) is introduced. Instead of increasing the prompt length, MoPE scales the parameter capacity of prompt fusion by adding more experts. This approach enhances expressiveness while avoiding the negative side effects associated with overly long prompt.
@@
We conduct systematic experiments on a total of six datasets spanning four modalities. Compared to existing prompt fusion methods, MoPE not only has better performance but also achieves a higher parameter efficiency. In our ablation analysis, we reveal that increasing the number of experts (i.e., “expert-scaling”) is more scalable than length-scaling, with monotonic performance gains and the avoidance of performance deterioration with an overly-long prompt. Furthermore, by introducing regularization terms for expert routing, we observe the emergence of specialized prompt experts after end-to-end training, resulting in high interpretability. Our key contributions are summarized as follows:
  • We propose a novel instance-wise adaptive prompt decomposition to augment the adaptiveness of prompt fusion.
@@
  • We introduce the MoPE technique for instance-wise dynamic prompt generation, which scales up the expressiveness of vanilla prompt.
@@
  • Extensive experiments on six datasets spanning four modalities demonstrate SOTA performance and parameter efficiency for prompt-based multimodal fusion.
2 Related Works
Prompt Tuning for Transfer Learning. Prompt tuning [19, 18] learns continuous token embeddings as additional input to a frozen pretrained model for transfer learning of Transformer-based models. It is widely used in various modalities [11, 3, 23, 22, 12, 14, 42]. A common observation is good transfer learning performance in low-shot scenarios, yet its performance is less comparable to fine-tuning when abundant training instances are available [19, 9]. Moreover, increasing prompt length quickly reaches performance saturation, and over-length prompts might lead to worse results [40, 11, 16]. Recent theoretical analyses [28, 37] reveal that the expressiveness of prompt tuning is lower than that of fine-tuning. In this work, we tackle this challenge by scaling up the expressiveness of prompts with a Mixture of Experts (MoE)-like design.
Prompt Fusion. Apart from model adaptation, prompts can also be used to fuse separately pretrained unimodal models for multimodal tasks. Frozen [34] first introduced a method where the visual representation is mapped as a few input tokens to query frozen language models (LMs). PromptFuse and BlindPrompt [21] improved upon this by introducing tunable prompts to the LM for cross-modal alignment. PICa [41] translates images into discrete text captions to prompt frozen LMs. These methods treat tokens from different modalities independently and lack explicit cross-modal interaction. Recently, PMF [20] introduced an interactive prompt fusion method for the vision and text modalities, building on the strong assumption that both encoders are white-box and have the same Transformer architecture. In this paper, we investigate how to adaptively interact with two or more modalities with minimal assumptions about the model architecture.
MoE in Transformers. MoE is a computationally efficient technique to scale up models, including Transformers [17, 31, 27, 8]. The fundamental approach involves inserting MoE layers, usually composed of multiple feed-forward networks (FFNs) acting as experts, into the standard Transformer architecture. A router is learned to route each token embedding to the most suitable expert(s) for reducing computational overhead [29, 31, 33]. In this paper, we do not focus on reducing computational cost; instead, we introduce the MoE design into prompt fusion to scale up its adaptiveness and expressiveness.
3 Method
@@
3.1 Preliminary: Vanilla Prompt Tuning
Consider a Transformer [35] or its variants used to extract features from an embedded input sequence $\mathbf{x}^{0}\in\mathbb{R}_{s\times d_{x}}$, where $s$ is the sequence length and $d_{x}$ is the embedding dimension. Prompt-tuning freezes all pre-trained parameters and optimize a small number of continuous embeddings (i.e., prompts) $\mathbf{P}\in\mathbb{R}_{l\times d}$ concatenated to the input of each layer, where $l$ is the length of the prompt. The input of the $i$-th layer layer $L^{i}$ could be denoted as:
@@ 3.1 Preliminary: Vanilla Prompt Tuning
Succinctly, prompt tuning works by biasing the pretrained attention pattern in each Transformer layers [28]. As opposed to fine-tuning, this biasing has strictly limited expressiveness in theory [28, 37], meaning that there are tasks that are un-learnable even with $l\to\infty$. This limitation characterizes the upper-bound of prompt tuning. On the other hand, the empirical expressiveness of vanilla prompting is even lower, which means that in practice, the prompt tuning tends to perform worse than its theoretical upper bound. This is due to competing optimization scheme when a long prompt $l>1$ [28] is employed. In other words, although longer prompt gives more trainable parameters, finding such a long prompt becomes challenging, and a suboptimal long prompt can lead to performance degradation instead. In this paper, we do not aim to increase the theoretical upper bound of prompt-tuning; instead, we approach the problem by narrowing the gap between its empirical and theoretical expressiveness.
3.2 Instance-wise Adaptive Prompt Decomposition
Previous prompt fusion methods [21, 41] optimize a global prompt that is shared across all instances, neglecting the interplay between different modalities and instance-specific features. With paired multimodal input as a prior, we aim to instance-wisely condition the prompting of one modality on the other(s) for better adaptivity. To achieve this goal, we first adopt a sequential fusion pipeline. Let $x\in\mathbb{X},y\in\mathbb{Y}$ be a pair of multimodal inputs, and $\mathcal{E}_{\mathbb{X}}$, $\mathcal{E}_{\mathbb{Y}}$ be the encoder of each modality. Depending on the task intrinsic, we assign a fusion direction where the encoding of main modality $\mathbb{X}$ is conditioned on representation of complementary modality(ies) $\mathbb{Y}$.
@@ 3.2 Instance-wise Adaptive Prompt Decomposition
3.3 Mixture of Prompt Experts
@@
where $\tau=0.1$ is the temperature hyper-parameter, and $\epsilon$ is sampled noise [31]. When there are more than one complementary modality, the additional embedding could be easily extended by learning additional projections. The dynamic prompt is obtained by a convex combination of all experts at this layer according to the routing score:
@@
We visualize the MoPE routing mechanism in Fig. 2-(b, c).
3.4 Regularizing Expert Routing
The proposed MoPE scales up vanilla prompt tuning by dividing the problem space into subspaces governed by specialized experts. However, we empirically find that a few experts tend to dominate across all instances during training, a phenomenon that has also been observed in previous MoE methods [7, 33]. In this section, we introduce methods to circumvent degenerated routing and aid expert specialization.
Orthogonal Routing Embedding. To avoid the expert being selected in initialization being exposed to a larger gradient signal, we freeze the routing embeddings $\{\mathbf{k}\}_{j=1}^{k}$ to suppress self-reinforcing expert domination. Our finding indicate that an orthogonal initialized [32] and non-learnable routing embedding outperform learned one, while requiring fewer trainable parameters.
Importance Loss. To further aid expert specialization, we add an additional importance loss [33, 31] to penalize dominant experts. For a batch of input $\mathbf{B}=\{(x_{0},y_{0}),(x_{1},y_{1}),...,(x_{b},y_{b})\}$, the importance of expert-$j$ is defined as the summed routing score in this batch, $\operatorname{Imp}(\mathbf{E}_{j})=\sum_{(x,y)\in\mathbf{B}}\mathbf{r}_{j}$. The importance loss is calculated as the mean coefficient of variation of all experts’ importance averaged across all layers:
@@ 3.4 Regularizing Expert Routing
where $\operatorname{\sigma}(\cdot)$ is the stop-gradient operator to prevent error propagation of this loss term when the coefficient of variation is less than a pre-defined threshold $\gamma=0.1$. While this loss was initially proposed for balancing computational budgets [33, 31], we adapt its use for promoting expert specialization. We add an additional threshold constraint due to our instance-wise routing, which is more likely to have a larger coefficient of variation than per-token routing.
@@
Acc % ($\uparrow$) Acc %($\uparrow$) F1-macro/F1-micro ($\uparrow$) Pre % ($\uparrow$) Pre % ($\uparrow$)
@@
4 Experiments
4.1 Dataset and Tasks
UPMC Food-101 [36] serves as a comprehensive multimodal dataset designed for fine-grained recipe classification. The dataset contains 90,840 image-text pairs for 101 food classes. We follow previous methods [20, 15] to create a validation set of 5000 samples.
MM-IMDB [1] is a multimodal movie classification dataset. It comprises 25,956 pairs of images and texts, each pair including a movie poster and a plot summary. The dataset supports multi-label genre classification across a spectrum of 23 genres with highly imbalanced classes.
SNLI-VE [39] is a large-scale multimodal dataset with 565,286 image-text pairs. The task for this dataset is visual entailment, which means that the model should decide whether a hypothesis matches the given premise or not. This dataset provide image and text premise, while the hypothesis is always in text modality.
MUStARD [4] is a dataset for multimodal sarcasm detection. The original dataset contains 690 video clips in MP4 format with an even number of sarcastic and non-sarcastic labels. Each video is provided with annotations in the language modality, describing the speaker and sentence in that video. Collectively, there are three modalities available: video, audio extracted from the video, and the language modality.
RefCOCO and RefCOCO+ [13] are two datasets for referring image segmentation. Both datasets are built based on MSCOCO and contain around 19,900 images with 50,000 objects and 140,000 referring expressions. RefCOCO allows referring expressions of any type, while RefCOCO+ features expressions that do not contain object positions.
4.2 Experiment Setups
Metrics. The metric on SNLI-VE, UPMC Food-101 is accuracy (%), MM-IMDB is F1-Macro and F1-Micro, and on MuSTARD is precision (%). Mean Intersection over Union (mIoU) is the metric for RefCOCO and RefCOCO+.
Architecture Details. In main experiment, we use pretrained Swin-B [25] as the image encoder, pretrained Bert-base [5] as the text encoder, and pretrained Wav2Vec2 [2] as the audio encoder. For video encoding, we follow PromptFuse [21] to use a ViT [6] for extracting image features from $n=8$ sampled frames, the averaged feature is used to represent the whole video.
Image is used as main modality unless otherwise specified. Following the experiment setup in [11, 20, 21], we finetune dataset-specific head. Linear head is used for all classification tasks, and we use a standard UperNet [38] head for segmentation task. We implement the mapper as a two-layer MLP with GeLU nonlinearity. Regarding the prompt, we set $l=6$ and use $k=16$ experts by default, which strike a balance between performance and parameter size. The prompts are applied to all layers of the main modality encoder. Vanilla prompt tuning [19] with $l^{\prime}=4$ is used to tune the $\mathcal{E}_{\mathbb{Y}}$. Further implementation details could be found in the Supplementary Material.
Training Details. All models are trained for 20 epoch, using the AdamW [26] optimizer with a learning rate of $4\times 10^{-4}$ for main modality and $5\times 10^{-4}$ for complementary modality. All models are trained with a RTX-4090 GPU.